# Few-Shot Adversarial Low-Rank Fine-Tuning of Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have shown remarkable performance in cross-modal tasks through large-scale contrastive pre-training. To adapt these large transformer-based models efficiently for downstream tasks, Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA have emerged as scalable alternatives to full fine-tuning, especially in few-shot scenarios. However, like traditional deep neural networks, VLMs are highly vulnerable to adversarial attacks, where imperceptible perturbations can significantly degrade model performance. Adversarial training remains the most effective strategy for improving model robustness in PEFT. In this work, we propose AdvCLIP-LoRA, the first algorithm designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings. Our method formulates adversarial fine-tuning as a minimax optimization problem and provides theoretical guarantees for convergence under smoothness and nonconvex-strong-concavity assumptions. Empirical results across eight datasets using ViT-B/16 and ViT-B/32 models show that AdvCLIP-LoRA significantly improves robustness against common adversarial attacks (e.g., FGSM, PGD), without sacrificing much clean accuracy. These findings highlight AdvCLIP-LoRA as a practical and theoretically grounded approach for robust adaptation of VLMs in resource-constrained settings.
@@ 2.2 Few-Shot Fine-Tuning for VLMs
To further adapt vision-language models to downstream tasks, the few-shot setting assumes access to a limited number of labeled examples per target class—typically fewer than $16$. Given $N$ such labeled support images per class, we denote the one-hot encoded ground-truth label for image $\mathbf{x}_{i}$ as $y_{ik}$, where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 00 otherwise. Classification probabilities $p_{i,k}$ are obtained as in the zero-shot setup, and the model is adapted by minimizing the cross-entropy loss: $\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$
@@ 3 Proposed Algorithm
5:    $\delta_{t}=\mathcal{P}_{\Delta}\left(\delta_{t-1}+\eta_{\delta}(\frac{1}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i}))\right)$triangleright Update and project the perturbation.
@@
7:    $B_{t}=B_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{B}F(W_{t-1},\delta_{t};\xi_{i})\right)$triangleright Update the low-rank matrices $A$ and $B$.
@@ 5 Empirical Results
5.1 Experiments Setup
Figure 3: Comparative analysis of CLIP-LoRA and AdvCLIP-LoRA with ViT-B/16 and ViT-B/32 backbones on $8$ fine-grained datasets, showing clean accuracy and PGD-adversarial robustness (shots labeled above). AdvCLIP-LoRAi means AdvCLIP-LoRA with $\tau=i$.
Adversarial Attacks. For adversarial training, we define the projection set for updating $\delta$ as an $\ell_{\infty}$-ball with a radius of $\epsilon=10$ across all datasets. To evaluate adversarial robustness, we implement two standard attack methods: FGSM [21] and PGD [23]. For FGSM, we set $\epsilon=10$, while for PGD, we use $\epsilon=2$ with a total of 20 attack iterations. Following the setup in [23], the step size for PGD is set to $\alpha=\frac{2.5\cdot\epsilon}{\texttt{Number of Iterations}}$.

### v2

@@ Abstract
Vision-Language Models (VLMs) such as CLIP have shown remarkable performance in cross-modal tasks through large-scale contrastive pre-training. To adapt these large transformer-based models efficiently for downstream tasks, Parameter-Efficient Fine-Tuning (PEFT) techniques like LoRA have emerged as scalable alternatives to full fine-tuning, especially in few-shot scenarios. However, like traditional deep neural networks, VLMs are highly vulnerable to adversarial attacks, where imperceptible perturbations can significantly degrade model performance. Adversarial training remains the most effective strategy for improving model robustness in PEFT. In this work, we propose AdvCLIP-LoRA, the first algorithm designed to enhance the adversarial robustness of CLIP models fine-tuned with LoRA in few-shot settings. Our method formulates adversarial fine-tuning as a minimax optimization problem and provides theoretical guarantees for convergence under smoothness and nonconvex-strong-concavity assumptions. Empirical results across eight datasets using ViT-B/16 and ViT-B/32 models show that AdvCLIP-LoRA significantly improves robustness against common adversarial attacks (e.g., FGSM, PGD), without sacrificing much clean accuracy. These findings highlight AdvCLIP-LoRA as a practical and theoretically grounded approach for robust adaptation of VLMs in resource-constrained settings. The code is available at https://github.com/sajjad-ucsb/AdvCLIP-LoRA.
@@ 2.2 Few-Shot Fine-Tuning for VLMs
To further adapt vision-language models to downstream tasks, the few-shot setting assumes access to a limited number of labeled examples per target class—typically fewer than $16$. Given $N$ such labeled support images per class, we denote the one-hot encoded ground-truth label for image $\mathbf{x}_{i}$ as $y_{ik}$, where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise. Classification probabilities $p_{i,k}$ are obtained as in the zero-shot setup, and the model is adapted by minimizing the cross-entropy loss: $\mathcal{L}_{\text{CE}}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$
@@ 3 Proposed Algorithm
5:  $\delta_{t}=\mathcal{P}_{\Delta}\left(\delta_{t-1}+\eta_{\delta}(\frac{1}{M}\sum_{i=1}^{M}\nabla_{\delta}F(W_{t-1},\delta_{t-1};\xi_{i}))\right)$ $\triangleright$ Update and project the perturbation.
@@
7:  $B_{t}=B_{t-1}-\eta_{w}\left(\frac{1}{M}\sum_{i=1}^{M}\nabla_{B}F(W_{t-1},\delta_{t};\xi_{i})\right)$ $\triangleright$ Update the low-rank matrices $A$ and $B$.
@@ 5 Empirical Results
Table 1: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold.
Average ImageNet Caltech DTD Food Pets Flowers UCF SUN
Shots Method Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD
1 C-AVP [37] 43.62 17.94 46.60 11.07 85.73 50.33 26.97 12.93 24.43 5.23 57.60 22.73 63.10 29.70 3.37 0.40 41.20 11.10
APT [24] 59.07 5.08 49.30 1.30 84.77 26.90 41.67 3.83 56.57 0.83 70.23 0.60 61.97 2.10 54.50 3.87 53.53 1.23
AdvPT [25] 29.96 1.44 20.17 0.43 62.97 7.60 16.73 2.60 13.27 0.00 37.93 0.13 33.97 0.43 27.03 0.00 27.57 0.37
AdvMaPLe [30] 33.52 11.38 49.27 14.60 85.53 48.37 13.63 2.93 5.27 0.30 30.67 4.97 1.40 0.10 32.70 7.07 49.70 12.67
AdvVLP [34] 32.82 11.78 49.00 15.53 85.43 48.47 15.97 4.77 1.07 0.77 29.63 3.83 19.77 6.57 11.83 1.73 49.83 12.60
FAP [34] 40.14 10.22 49.90 15.40 83.53 41.13 18.40 2.40 31.67 1.43 49.23 3.47 10.40 0.53 28.50 2.43 49.53 14.93
AdvCLIP-LoRA1 67.18 23.9 50.86 15.89 90.79 59.31 47.70 14.24 74.22 11.81 85.80 29.14 64.27 18.64 63.73 17.31 60.05 24.90
AdvCLIP-LoRA2 66.12 24.24 51.94 17.19 87.42 58.22 48.35 15.31 73.34 11.94 83.18 29.30 63.74 19.04 61.70 16.92 59.30 25.99
AdvCLIP-LoRA5 64.22 26.32 50.29 17.53 85.56 60.93 47.75 16.96 72.82 15.01 79.39 32.92 60.94 19.57 58.90 19.83 58.10 27.78
AdvCLIP-LoRA10 61.70 27.12 47.57 17.66 83.61 60.41 47.22 16.43 69.66 18.10 74.93 32.71 60.13 22.05 53.85 21.02 56.62 28.61
AdvCLIP-LoRA25 53.32 26.28 36.96 14.74 79.07 61.46 45.74 18.79 55.32 17.59 62.77 30.36 49.86 20.30 45.33 19.67 51.52 27.29
AdvCLIP-LoRA50 34.78 19.41 27.66 10.44 68.15 57.69 39.83 18.91 17.17 4.47 29.38 14.80 15.14 8.57 36.06 17.71 44.88 22.71
AdvCLIP-LoRA75 27.75 16.53 22.18 8.26 64.54 54.52 36.23 20.15 10.20 2.67 17.66 9.27 3.74 2.80 29.08 14.17 38.39 20.43
AdvCLIP-LoRA100 23.23 14.81 15.51 5.57 62.27 52.86 30.61 19.50 8.15 2.17 10.00 6.57 3.21 3.13 25.27 12.61 30.80 16.08
4 C-AVP [37] 43.10 16.40 49.80 11.13 90.17 52.50 18.77 9.27 22.73 4.57 57.80 16.20 55.97 23.73 1.07 0.80 48.47 13.03
APT [24] 66.37 6.04 50.90 1.40 90.77 26.67 51.33 6.33 54.80 1.63 71.83 2.10 82.40 4.23 66.53 3.03 62.37 2.90
AdvPT [25] 35.32 2.07 23.40 1.33 64.97 7.30 31.70 4.37 15.23 0.37 44.13 1.73 41.97 0.63 31.17 0.47 29.97 0.40
AdvMaPLe [30] 51.01 21.61 51.27 19.00 89.53 59.40 6.43 2.40 60.00 14.83 30.70 9.03 52.20 25.37 59.73 21.30 58.23 21.53
AdvVLP [34] 55.18 23.40 51.30 19.37 89.37 59.07 22.97 10.33 41.50 11.20 67.43 18.47 51.00 25.80 59.97 21.77 57.90 21.17
FAP [34] 57.51 24.6 51.53 19.60 87.57 57.33 31.27 8.07 59.37 18.37 42.10 9.30 73.13 38.77 58.50 22.13 56.60 23.20
AdvCLIP-LoRA1 76.11 31.89 62.53 21.89 93.59 69.41 58.69 21.87 77.21 15.79 86.86 35.65 86.20 35.16 74.20 24.11 69.64 31.26
AdvCLIP-LoRA2 76.24 34.05 62.47 22.91 93.35 71.40 59.69 23.88 77.29 18.10 87.68 37.48 85.91 39.26 73.91 26.51 69.60 32.84
AdvCLIP-LoRA5 75.75 37.12 62.11 24.69 93.10 73.59 57.51 27.13 76.49 20.91 88.12 40.58 85.79 45.11 73.80 29.58 69.09 35.39
AdvCLIP-LoRA10 75.09 38.07 61.41 25.43 93.27 74.20 57.57 27.54 75.11 23.01 87.54 41.40 85.51 48.40 71.82 28.26 68.49 36.29
AdvCLIP-LoRA25 72.93 39.26 58.08 25.56 93.18 75.90 53.90 27.84 73.33 26.25 87.11 41.26 81.24 50.51 70.53 30.82 66.03 36.94
AdvCLIP-LoRA35 71.52 39.12 56.40 24.58 93.10 78.30 53.07 26.89 71.65 25.47 85.88 41.16 79.54 48.84 67.88 31.03 64.66 36.68
AdvCLIP-LoRA50 69.21 38.42 54.52 23.33 92.41 76.84 51.65 26.65 68.55 24.57 83.95 39.98 75.23 48.40 64.34 31.96 63.03 35.65
AdvCLIP-LoRA75 63.54 35.21 46.98 19.83 91.20 75.70 48.17 26.71 57.67 19.82 82.77 37.53 64.35 40.48 57.71 28.36 59.46 33.23
AdvCLIP-LoRA100 52.08 30.12 20.24 10.42 88.64 73.59 44.98 24.70 22.08 7.55 81.22 35.57 53.35 33.37 53.56 26.43 52.54 29.33
16 C-AVP [37] 41.90 17.13 46.27 12.77 90.40 52.60 29.20 13.87 1.07 0.80 56.40 16.43 56.17 22.03 0.97 0.93 54.70 17.63
APT [24] 71.05 8.35 52.63 2.07 92.93 30.23 54.93 10.47 62.50 2.63 83.70 4.40 86.63 8.97 69.40 4.40 65.67 3.67
AdvPT [25] 40.94 2.68 24.53 1.47 68.70 9.63 43.77 5.70 18.47 0.73 46.27 0.23 56.03 0.80 36.60 0.53 33.13 2.37
AdvMaPLe [30] 71.48 38.11 52.93 21.90 92.17 68.63 57.93 32.17 65.13 25.27 83.27 36.87 87.87 58.70 68.97 31.67 63.57 29.70
AdvVLP [34] 68.76 37.01 53.23 22.10 92.37 67.97 57.53 32.73 43.30 16.50 82.93 35.57 87.70 58.70 69.10 32.80 63.90 29.70
FAP [34] 69.88 39.22 52.53 22.90 91.10 67.33 55.17 31.33 64.03 26.67 81.90 41.00 86.27 61.47 65.70 32.80 62.37 30.27
AdvCLIP-LoRA1 80.91 35.09 66.83 24.88 95.29 72.66 66.55 25.30 78.68 17.07 88.74 33.17 95.62 48.56 81.68 25.17 73.87 33.88
AdvCLIP-LoRA2 80.96 38.25 66.74 26.03 95.25 75.38 66.08 29.61 78.52 19.71 89.21 36.55 95.82 52.82 82.10 28.60 73.99 37.26
AdvCLIP-LoRA5 80.61 40.24 66.59 28.11 95.74 76.67 65.37 30.20 78.36 22.14 89.13 37.56 95.86 56.03 79.99 32.09 73.84 39.15
AdvCLIP-LoRA10 80.12 41.53 66.06 29.48 95.38 76.80 64.60 31.74 77.57 24.21 88.66 37.53 95.53 57.53 79.70 34.58 73.49 40.35
AdvCLIP-LoRA25 78.73 42.40 64.52 30.50 95.42 78.26 63.00 31.86 75.34 24.50 86.86 38.38 94.32 58.95 77.82 34.63 72.55 42.14
AdvCLIP-LoRA35 77.49 41.97 63.06 29.81 95.21 78.78 61.88 30.97 73.74 23.61 86.07 37.88 93.10 59.11 75.15 33.70 71.70 41.89
AdvCLIP-LoRA50 75.74 39.24 61.45 28.74 94.81 78.95 58.87 29.91 70.75 20.81 85.01 35.60 92.12 50.99 72.61 27.73 70.27 40.17
AdvCLIP-LoRA75 72.41 35.04 58.38 25.86 93.55 74.73 55.08 26.77 66.56 18.21 82.15 29.93 89.69 47.06 69.36 23.10 64.54 34.63
AdvCLIP-LoRA100 68.39 31.50 46.58 21.85 91.72 71.60 53.43 25.06 61.82 13.60 80.19 25.29 86.85 40.32 65.13 22.10 61.40 31.17
5.1 Experiments Setup
Adversarial Attacks. For adversarial training, we define the projection set for updating $\delta$ as an $\ell_{\infty}$-ball with a radius of $\epsilon\in\{1/255,10/255\}$ across all datasets. To evaluate adversarial robustness, we implement two standard attack methods: FGSM [21] and PGD [23]. For FGSM, we set $\epsilon=10/255$, while for PGD, we use $\epsilon\in\{1/255,2/255\}$ with a total of 20 attack iterations. Following the setup in [23], the step size for PGD is set to $\alpha=\frac{2.5\cdot\epsilon}{\texttt{Number of Iterations}}$.
5.2 Comparative Analysis of AdvCLIP-LoRA with Robust PEFT Methods
We conduct a thorough comparison between AdvCLIP-LoRA and competitive baselines using ViT-B/32 as the backbone across 8 datasets. Table 1 presents the results for 1-, 4-, and 16-shot settings. Complete results and corresponding plots are provided in Table 3 and Fig. B in the Appendix. Our baselines include state-of-the-art methods such as C-AVP [1], APT [2], AdvPT [3], AdvMaPLe [4], AdvVLP [5], and FAP [5]. We follow the setup in FAP [5] for measuring adversarial robustness using a 2-step PGD attack with an $\ell_{\infty}$ norm, a perturbation bound of $\epsilon=1/255$, and a step size of $\alpha=1/255$. For a fair comparison, we report results for $\tau=1,2$ (matching the baselines) and also explore higher $\tau$ values to illustrate the enhanced adversarial fine-tuning capabilities of our method.
On average, over all datasets:
  • Clean Accuracy: AdvCLIP-LoRA with $\tau=2$ outperforms the best baseline (FAP) by 25.98, 16.66, 18.75, 11.35, and 11.08 percentage points for 1, 2, 4, 8, and 16 shots, respectively.
  • Robust Accuracy: AdvCLIP-LoRA with $\tau=2$ exceeds FAP by 14.02, 11.71, 9.45, and 0.87 percentage points for 1, 2, 4, and 8 shots; for 16 shots, it is only 0.97 percentage points below FAP.
We also note that in our method, increasing $\tau$ up to a moderate threshold (about 25 ) can further improve the clean/robust trade-off for most tasks, whereas the effect on other baselines is unclear and would require more extensive retuning.
5.3 Practical Guidance on Selecting the Hyperparameter $\tau$
According to Figure 1 (full version in Fig 3 in the Appendix), our observations on the choice of $\tau$ are as follows:
  • Optimal Range for Robustness: For most tasks, robust accuracy continues to improve up to $\tau=25$ for few-shot settings ( $2,4,8$, and 16 shots) and up to $\tau=10$ for the 1-shot setting.
  • Point of Diminishing Returns: Beyond $\tau=25$, we observe that for the majority of datasets, both clean and robust accuracy begin to decrease. This suggests the model may start to overfit to the specific adversarial examples generated during training, harming its generalizability.
  • Dataset-Specific Behavior: We note some tasks, like Caltech101, can benefit from even larger $\tau$ values (up to 35 ), highlighting that the optimal $\tau$ can have some data dependency.
  • Clear Trade-off: The results confirm a clear trade-off where larger $\tau$ values consistently improve robustness up to a point, but at a predictable cost to clean accuracy.
From this analysis, we recommend setting $\tau$ between 10 and 25 as a practical choice for achieving a strong balance between clean and robust accuracy in few-shot learning.
@@ 5.3 Practical Guidance on Selecting the Hyperparameter $\tau$
Figure 3: Comparative analysis of CLIP-LoRA and AdvCLIP-LoRA with ViT-B/16 and ViT-B/32 backbones on 88 fine-grained datasets, showing clean accuracy and PGD-adversarial robustness (shots labeled above). AdvCLIP-LoRAi means AdvCLIP-LoRA with $\tau=i$.
@@ Appendix A Experiment Details
Table 3: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold.
Average ImageNet Caltech DTD Food Pets Flowers UCF SUN
Shots Method Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD Clean PGD
1 C-AVP [37] 43.62 17.94 46.60 11.07 85.73 50.33 26.97 12.93 24.43 5.23 57.60 22.73 63.10 29.70 3.37 0.40 41.20 11.10
APT [24] 59.07 5.08 49.30 1.30 84.77 26.90 41.67 3.83 56.57 0.83 70.23 0.60 61.97 2.10 54.50 3.87 53.53 1.23
AdvPT [25] 29.96 1.44 20.17 0.43 62.97 7.60 16.73 2.60 13.27 0.00 37.93 0.13 33.97 0.43 27.03 0.00 27.57 0.37
AdvMaPLe [30] 33.52 11.38 49.27 14.60 85.53 48.37 13.63 2.93 5.27 0.30 30.67 4.97 1.40 0.10 32.70 7.07 49.70 12.67
AdvVLP [34] 32.82 11.78 49.00 15.53 85.43 48.47 15.97 4.77 1.07 0.77 29.63 3.83 19.77 6.57 11.83 1.73 49.83 12.60
FAP [34] 40.14 10.22 49.90 15.40 83.53 41.13 18.40 2.40 31.67 1.43 49.23 3.47 10.40 0.53 28.50 2.43 49.53 14.93
AdvCLIP-LoRA1 67.18 23.9 50.86 15.89 90.79 59.31 47.70 14.24 74.22 11.81 85.80 29.14 64.27 18.64 63.73 17.31 60.05 24.90
AdvCLIP-LoRA2 66.12 24.24 51.94 17.19 87.42 58.22 48.35 15.31 73.34 11.94 83.18 29.30 63.74 19.04 61.70 16.92 59.30 25.99
AdvCLIP-LoRA5 64.22 26.32 50.29 17.53 85.56 60.93 47.75 16.96 72.82 15.01 79.39 32.92 60.94 19.57 58.90 19.83 58.10 27.78
AdvCLIP-LoRA10 61.70 27.12 47.57 17.66 83.61 60.41 47.22 16.43 69.66 18.10 74.93 32.71 60.13 22.05 53.85 21.02 56.62 28.61
AdvCLIP-LoRA25 53.32 26.28 36.96 14.74 79.07 61.46 45.74 18.79 55.32 17.59 62.77 30.36 49.86 20.30 45.33 19.67 51.52 27.29
AdvCLIP-LoRA50 34.78 19.41 27.66 10.44 68.15 57.69 39.83 18.91 17.17 4.47 29.38 14.80 15.14 8.57 36.06 17.71 44.88 22.71
AdvCLIP-LoRA75 27.75 16.53 22.18 8.26 64.54 54.52 36.23 20.15 10.20 2.67 17.66 9.27 3.74 2.80 29.08 14.17 38.39 20.43
AdvCLIP-LoRA100 23.23 14.81 15.51 5.57 62.27 52.86 30.61 19.50 8.15 2.17 10.00 6.57 3.21 3.13 25.27 12.61 30.80 16.08
2 C-AVP [37] 39.24 16.23 46.23 10.90 91.25 55.23 14.27 6.93 1.05 0.10 47.13 15.10 61.47 26.93 1.73 1.07 50.77 13.57
APT [24] 63.56 5.90 48.83 1.03 89.70 31.70 45.57 4.27 60.17 0.87 72.87 1.07 67.17 3.10 65.00 3.10 59.20 2.03
AdvPT [25] 32.47 1.76 22.37 0.77 66.07 8.33 24.57 2.43 11.13 0.17 39.17 1.17 38.47 0.23 29.37 0.23 28.57 0.77
AdvMaPLe [30] 39.09 15.58 49.97 17.13 88.00 56.20 16.53 4.20 3.10 0.67 34.03 6.87 46.17 17.00 21.17 6.20 53.73 16.33
AdvVLP [34] 42.79 17.76 50.53 17.50 87.60 55.33 18.33 7.17 1.53 1.10 31.27 7.07 62.43 25.17 36.83 11.43 53.77 17.33
FAP [34] 55.18 18.14 48.53 17.83 87.73 53.90 18.40 4.33 56.90 10.53 64.23 12.67 53.10 19.57 58.50 7.03 54.07 19.30
AdvCLIP-LoRA1 72.09 28.02 59.25 20.41 91.72 64.22 53.55 19.86 76.00 15.23 84.30 29.65 78.56 25.09 67.86 21.52 65.51 28.15
AdvCLIP-LoRA2 71.84 29.85 59.27 20.88 91.32 66.17 52.78 20.57 76.27 17.05 83.57 32.95 78.20 29.56 67.62 22.05 65.73 29.58
AdvCLIP-LoRA5 71.70 33.6 58.44 21.91 91.20 68.80 53.07 22.93 75.91 21.15 84.49 39.71 77.47 35.44 67.86 26.88 65.13 31.94
AdvCLIP-LoRA10 71.15 34.8 57.44 22.92 90.99 71.16 52.96 24.29 74.93 21.99 83.48 38.62 76.74 37.72 68.57 28.05 64.07 33.68
AdvCLIP-LoRA25 68.53 36.47 52.66 21.90 90.83 73.87 52.60 28.25 72.09 22.42 83.59 43.96 71.42 40.52 64.39 27.33 60.69 33.51
AdvCLIP-LoRA35 65.92 35.82 49.24 20.08 88.60 73.79 51.54 26.83 69.83 22.63 81.82 42.85 68.01 39.99 58.71 28.07 58.63 32.31
AdvCLIP-LoRA50 61.09 33.98 42.74 16.95 86.41 73.02 49.05 26.48 57.83 18.46 80.46 40.53 62.08 38.45 53.77 26.91 56.34 31.01
AdvCLIP-LoRA75 52.39 30.60 32.63 13.46 82.68 69.82 46.99 27.72 36.20 10.63 72.20 35.62 48.36 32.97 48.61 25.17 52.45 29.44
AdvCLIP-LoRA100 37.08 24.24 16.63 7.92 80.61 67.71 45.45 29.79 12.30 3.96 38.35 23.28 19.33 14.13 42.37 22.79 41.59 24.31
4 C-AVP [37] 43.10 16.40 49.80 11.13 90.17 52.50 18.77 9.27 22.73 4.57 57.80 16.20 55.97 23.73 1.07 0.80 48.47 13.03
APT [24] 66.37 6.04 50.90 1.40 90.77 26.67 51.33 6.33 54.80 1.63 71.83 2.10 82.40 4.23 66.53 3.03 62.37 2.90
AdvPT [25] 35.32 2.07 23.40 1.33 64.97 7.30 31.70 4.37 15.23 0.37 44.13 1.73 41.97 0.63 31.17 0.47 29.97 0.40
AdvMaPLe [30] 51.01 21.61 51.27 19.00 89.53 59.40 6.43 2.40 60.00 14.83 30.70 9.03 52.20 25.37 59.73 21.30 58.23 21.53
AdvVLP [34] 55.18 23.40 51.30 19.37 89.37 59.07 22.97 10.33 41.50 11.20 67.43 18.47 51.00 25.80 59.97 21.77 57.90 21.17
FAP [34] 57.51 24.6 51.53 19.60 87.57 57.33 31.27 8.07 59.37 18.37 42.10 9.30 73.13 38.77 58.50 22.13 56.60 23.20
AdvCLIP-LoRA1 76.11 31.89 62.53 21.89 93.59 69.41 58.69 21.87 77.21 15.79 86.86 35.65 86.20 35.16 74.20 24.11 69.64 31.26
AdvCLIP-LoRA2 76.24 34.05 62.47 22.91 93.35 71.40 59.69 23.88 77.29 18.10 87.68 37.48 85.91 39.26 73.91 26.51 69.60 32.84
AdvCLIP-LoRA5 75.75 37.12 62.11 24.69 93.10 73.59 57.51 27.13 76.49 20.91 88.12 40.58 85.79 45.11 73.80 29.58 69.09 35.39
AdvCLIP-LoRA10 75.09 38.07 61.41 25.43 93.27 74.20 57.57 27.54 75.11 23.01 87.54 41.40 85.51 48.40 71.82 28.26 68.49 36.29
AdvCLIP-LoRA25 72.93 39.26 58.08 25.56 93.18 75.90 53.90 27.84 73.33 26.25 87.11 41.26 81.24 50.51 70.53 30.82 66.03 36.94
AdvCLIP-LoRA35 71.52 39.12 56.40 24.58 93.10 78.30 53.07 26.89 71.65 25.47 85.88 41.16 79.54 48.84 67.88 31.03 64.66 36.68
AdvCLIP-LoRA50 69.21 38.42 54.52 23.33 92.41 76.84 51.65 26.65 68.55 24.57 83.95 39.98 75.23 48.40 64.34 31.96 63.03 35.65
AdvCLIP-LoRA75 63.54 35.21 46.98 19.83 91.20 75.70 48.17 26.71 57.67 19.82 82.77 37.53 64.35 40.48 57.71 28.36 59.46 33.23
AdvCLIP-LoRA100 52.08 30.12 20.24 10.42 88.64 73.59 44.98 24.70 22.08 7.55 81.22 35.57 53.35 33.37 53.56 26.43 52.54 29.33
8 C-AVP [37] 43.24 17.21 46.37 11.90 91.20 50.33 23.63 11.47 1.00 0.83 57.43 17.33 55.50 23.57 18.27 4.93 52.53 17.30
APT [24] 69.76 7.56 52.03 1.80 92.37 30.83 54.43 8.70 61.57 2.33 82.87 3.10 84.00 6.00 66.53 4.30 64.30 3.40
AdvPT [25] 38.50 2.39 24.30 0.87 68.07 10.10 37.47 4.20 16.97 0.10 44.20 0.27 51.13 0.87 33.43 1.40 32.47 1.33
AdvMaPLe [30] 62.90 30.45 52.13 20.60 90.63 63.80 33.20 16.97 62.70 20.13 55.60 21.07 83.10 48.80 64.33 25.93 61.50 26.30
AdvVLP [34] 68.32 32.87 52.83 20.97 90.17 63.13 51.83 25.77 61.73 19.33 80.67 29.63 83.90 50.90 64.07 26.97 61.33 26.23
FAP [34] 67.23 34.26 52.17 21.53 89.63 62.50 52.13 25.77 61.80 23.20 79.47 34.57 81.53 52.63 60.70 26.67 60.40 27.23
AdvCLIP-LoRA1 78.82 33.28 64.97 23.16 94.04 70.06 62.06 26.42 78.07 15.99 88.91 34.37 91.35 38.29 79.33 25.75 71.83 32.18
AdvCLIP-LoRA2 78.58 35.13 64.78 24.44 94.00 72.05 61.17 22.75 78.26 19.18 88.74 36.88 90.74 44.50 79.12 26.91 71.80 34.37
AdvCLIP-LoRA5 78.01 38.13 64.53 26.49 94.16 73.91 59.75 29.49 77.61 22.26 88.83 38.27 90.50 48.92 77.24 28.60 71.44 37.11
AdvCLIP-LoRA10 77.50 39.61 63.98 27.83 93.91 74.40 59.63 31.09 76.52 24.00 88.01 39.36 90.62 52.13 76.71 30.06 70.64 38.00
AdvCLIP-LoRA25 75.18 40.24 61.98 28.66 93.55 74.81 54.31 28.61 73.77 25.47 86.45 38.48 88.71 54.53 72.98 32.01 69.65 39.36
AdvCLIP-LoRA35 73.93 39.98 60.67 27.93 93.39 75.82 54.31 28.49 71.56 25.06 86.02 39.19 86.80 52.33 70.21 32.49 68.46 38.50
AdvCLIP-LoRA50 71.75 38.55 58.75 26.69 92.86 75.29 50.89 27.30 69.10 24.62 84.46 36.79 84.08 49.05 66.88 31.22 67.00 37.43
AdvCLIP-LoRA75 69.16 35.98 56.12 24.84 91.68 73.79 49.35 25.30 65.35 21.86 82.39 33.03 81.12 46.00 63.65 28.73 63.61 34.28
AdvCLIP-LoRA100 64.88 33.85 47.53 22.94 90.47 71.32 46.63 23.88 59.80 22.16 80.46 29.65 74.06 40.11 60.53 28.87 60.54 31.84
16 C-AVP [37] 41.90 17.13 46.27 12.77 90.40 52.60 29.20 13.87 1.07 0.80 56.40 16.43 56.17 22.03 0.97 0.93 54.70 17.63
APT [24] 71.05 8.35 52.63 2.07 92.93 30.23 54.93 10.47 62.50 2.63 83.70 4.40 86.63 8.97 69.40 4.40 65.67 3.67
AdvPT [25] 40.94 2.68 24.53 1.47 68.70 9.63 43.77 5.70 18.47 0.73 46.27 0.23 56.03 0.80 36.60 0.53 33.13 2.37
AdvMaPLe [30] 71.48 38.11 52.93 21.90 92.17 68.63 57.93 32.17 65.13 25.27 83.27 36.87 87.87 58.70 68.97 31.67 63.57 29.70
AdvVLP [34] 68.76 37.01 53.23 22.10 92.37 67.97 57.53 32.73 43.30 16.50 82.93 35.57 87.70 58.70 69.10 32.80 63.90 29.70
FAP [34] 69.88 39.22 52.53 22.90 91.10 67.33 55.17 31.33 64.03 26.67 81.90 41.00 86.27 61.47 65.70 32.80 62.37 30.27
AdvCLIP-LoRA1 80.91 35.09 66.83 24.88 95.29 72.66 66.55 25.30 78.68 17.07 88.74 33.17 95.62 48.56 81.68 25.17 73.87 33.88
AdvCLIP-LoRA2 80.96 38.25 66.74 26.03 95.25 75.38 66.08 29.61 78.52 19.71 89.21 36.55 95.82 52.82 82.10 28.60 73.99 37.26
AdvCLIP-LoRA5 80.61 40.24 66.59 28.11 95.74 76.67 65.37 30.20 78.36 22.14 89.13 37.56 95.86 56.03 79.99 32.09 73.84 39.15
AdvCLIP-LoRA10 80.12 41.53 66.06 29.48 95.38 76.80 64.60 31.74 77.57 24.21 88.66 37.53 95.53 57.53 79.70 34.58 73.49 40.35
AdvCLIP-LoRA25 78.73 42.40 64.52 30.50 95.42 78.26 63.00 31.86 75.34 24.50 86.86 38.38 94.32 58.95 77.82 34.63 72.55 42.14
AdvCLIP-LoRA35 77.49 41.97 63.06 29.81 95.21 78.78 61.88 30.97 73.74 23.61 86.07 37.88 93.10 59.11 75.15 33.70 71.70 41.89
AdvCLIP-LoRA50 75.74 39.24 61.45 28.74 94.81 78.95 58.87 29.91 70.75 20.81 85.01 35.60 92.12 50.99 72.61 27.73 70.27 40.17
AdvCLIP-LoRA75 72.41 35.04 58.38 25.86 93.55 74.73 55.08 26.77 66.56 18.21 82.15 29.93 89.69 47.06 69.36 23.10 64.54 34.63
AdvCLIP-LoRA100 68.39 31.50 46.58 21.85 91.72 71.60 53.43 25.06 61.82 13.60 80.19 25.29 86.85 40.32 65.13 22.10 61.40 31.17
Figure 7: Detailed comparative analysis of various adversarial PEFT methods with ViT-B/32 as backbone.